OpenStudy (angelina__nievse):

Simplify 9^2/ 9^7 will give a medal

OpenStudy (junebuggie12):

9^2 would be 9*9. 9^7 would be 9*9*9*9*9*9*9. Solve those for your answers.

OpenStudy (whpalmer4):

\[\frac{9^2}{9^7} = \frac{9*9}{9*9*9*9*9*9*9} = \] Some useful properties of exponents: \[\frac{b^m}{b^n} = b^{m-n}\] \[\frac{1}{b^m} = b^{-m}\]